# Chameleon: Automated Color Palette Adaptation for Dark Mode Data Visualizations

### v1

@@ Abstract.
Figure 1. Visualization examples adapted from Shen et al. (Shen2017NameClarifierAV), demonstrating four different rendering conditions. (A) Original visualization in light mode. (B) A typical dark mode conversion, where only the background is darkened while the original foreground colors are left unchanged, often leading to poor contrast and reduced legibility. (C) Visualization rendered using direct color inversion, which can distort the intended color semantics and lead to misinterpretation. (D) Visualization transformed using our method, Chameleon, which optimizes color contrast and preserves semantic color mappings for improved legibility in dark mode.
@@ 1. Introduction
In recent years, dark mode has gained popularity across mobile platforms like Android and iOS due to its benefits, such as reduced eye strain in low-light conditions and prolonged battery life on OLED and AMOLED screens (Andrew2024LightAD; Iyer2003EnergyAdaptiveDS). Dark mode is a user interface (UI) theme that uses a dark background with light text, as opposed to the default light mode, which typically uses a light background with dark text. Users can easily switch to dark mode through system settings, resulting in an immediate change in the system’s UI. Visualizations are commonly used on mobile devices because of their convenience (Choe2019MobileDV; Lee2020ReachingBA; Roberts2014VisualizationBT). When dark mode visualizations are effectively designed, they minimize distractions (appleDarkMode), make important content stand out, and reduce eye strain (Qiao2023LightMA; Popelka2022TheEO). They can also be made to adjust to changing light conditions, improving user comfort in dim environments (Qiao2023LightMA). For instance, Mairena et al. (Mairena2021WhichET) found that the users performed better using the dark mode scatterplots, with faster search time and higher accuracy, compared to light mode.
@@
However, most visualizations were designed for light mode, and when the system switches to dark mode, the mismatch between visualization elements and the dark background creates significant visibility problems. Figure 1[Uncaptioned image] shows a visualization from a scientific paper (Shen2017NameClarifierAV) that is clearly visible against a light background. Yet, when the system switches to dark mode, as shown in Figure 1[Uncaptioned image], the visualization elements become difficult to discern against the dark background, compromising their legibility. A common shortcut is to simply invert the color palette (Andrew2024LightAD), as shown in Figure 1[Uncaptioned image]. Unfortunately, this inversion cannot preserve the color semantic relationships between the inverted visualization and the original visualization: the blue color in the original visualization is inverted to yellow, and the orange color is inverted to blue. This change largely disrupts users’ cognitive ability to recognize and interpret the visualization across both modes.
@@
As a consequence, designers are left to manually adapt visualizations for dark mode. This approach requires designers to build dark mode visualizations from scratch, a time-consuming process that involves complex multi-factor considerations: maintaining adequate contrast against the dark background, preserving the original color relationships, and ensuring overall visual harmony. These factors often conflict with each other. For example, increasing contrast for better legibility might disrupt color relationships or visual harmony. Since such factors are interdependent, designers must engage in extensive trial and error, making repeated adjustments to balance these competing requirements. While operating systems like Android and iOS offer guidelines (appleColorApple; appleDarkMode; androidDarkMode; pinterest) for UI design in dark mode, these guidelines typically focus on discrete UI elements like buttons and menus. As a consequence, the guidelines are often too abstract to be directly applicable to the specific needs of visualizations (Andrew2024LightAD; Qiao2023LightMA), which requires more nuanced design considerations.
@@
To reduce designers’ workload, there is a need for an automated tool that can transform light mode visualizations to dark mode by adjusting their color palettes. While there are existing tools for automatic color palette generation (Lu2020PalettailorDC; Gramazio2017ColorgoricalCD; Nardini2021AutomaticIO), these cannot be directly applied to dark mode transformation. These tools focus solely on generating optimized palettes for single mode and do not address the challenges of cross-mode transformation. These single-mode approaches creates several problems. First, colors that appear legible in light mode may become harsh or difficult to distinguish against dark backgrounds, an issue not considered by existing automated methods. Second, colors need to maintain consistency between the modes to help users quickly interpret visualizations regardless of display mode. Finally, visual harmony and hierarchy must be preserved. For example, if highlighted regions have distinct contrast with adjacent areas in the light mode, this relationship should persist in the dark mode for consistent information emphasis. These cross-mode relationships and dark background interactions present unique challenges that require a different approach to the color palette generation.
@@
To fill this gap, we propose Chameleon, an algorithm designed to automatically transform visualizations from light mode to dark mode. Through literature reviews (appleDarkMode; androidDarkMode; Lu2020PalettailorDC; Gramazio2017ColorgoricalCD; Nardini2021AutomaticIO), we identified three critical factors in dark mode visualization design: maintaining adequate luminance contrast for legibility, preserving color semantic relationships, and ensuring visual harmony between adjacent colors. The key challenge lies in quantifying these factors and balancing them effectively. Chameleon addresses this by formulating the transformation as an optimization problem that mathematically models these design factors. Specifically, Chameleon takes an image of a light mode visualization as input and analyzes its color palette. The algorithm then optimizes the dark mode color palette by balancing three designed loss functions: luminance contrast consistency to ensure legibility against the dark background, color consistency to maintain semantic relationships from the original visualization, and adjacent color difference consistency to preserve visual harmony and hierarchy. This automated optimization eliminates the initial trial-and-error phase, providing designers with an optimized starting point that they can fine-tune rather than starting from scratch. The weights of these loss functions can be adjusted if designers want to prioritize different factors according to their specific needs. As shown in Figure 1[Uncaptioned image] and Figure 1[Uncaptioned image], this approach produces a dark mode visualization that successfully balances visual clarity, color consistency, and visual harmony.
@@ 2.1. Color Palette Design Principles
Color palette design for charts centers on two main considerations: perceptual discriminability and aesthetics (Yuan2021InfoColorizerIR). The perceptual discriminability ensures that users can easily distinguish between colors within a visualization (Gramazio2017ColorgoricalCD; Stone2014AnEM), a requirement that is critical for accurate data interpretation. Healey (Healey1996ChoosingEC) emphasized that “colors should be well-separated,” underscoring the importance of maximizing visual distinctness. Several factors influence discriminability: the data properties being represented (e.g., continuous vs. categorical data) (Levkowitz1992ColorSF; Tennekes2014TreeCC), the analytic tasks users perform (e.g., comparing classes, identifying outliers) (Tominski2008TaskDrivenCC), the ease of color naming and categorical labeling (Heer2012ColorNM), and the degree of perceptual contrast among colors (Wang2019OptimizingCA; Mittelstdt2014MethodsFC). Healey (Healey1996ChoosingEC) proposed selecting representative colors from ten hue regions in Munsell’s color space while maximizing perceptual distances between them, while Maxwell (maxwell2000visualizing) similarly sought to maximize class discriminability, accounting for spatial distribution of colors. However, these approaches face limitations like the resulting palettes are not always aesthetically pleasing (Lu2020PalettailorDC), stressing that a proper balance between perceptual discriminability and aesthetics.
Beyond discriminability, palettes must also be visually appealing, as aesthetic qualities influence user engagement and interpretability. Research shows that color harmony, the balanced arrangement of colors, enhances visual appeal and reduces cognitive strain (Meier2004InteractiveCP). Furthermore, manipulating hue, saturation, and brightness allows designers to balance vividness with readability, ensuring that palettes remain both expressive and coherent (Moreland2009DivergingCM). Well-chosen palettes thus balance clarity and beauty, encouraging users to focus on insights rather than being distracted by jarring color choices. These design principles take on added importance in dark mode settings, where background inversion can drastically alter perceived contrast and harmony. Guidelines from major platforms such as Android (androidDarkMode) and iOS (appleDarkMode) stress the importance of maintaining strong luminance contrast between visualization elements and background. Without such care, a palette optimized for light mode may become unreadable, misleading, or aesthetically unbalanced in dark mode environments. In this work, we build on these principles to address the unique challenges of dark mode visualization.
@@ 2.2. Color Palette Generation
The existing approaches to color palette generation fall into two categories: pre-defined and automatic. Pre-defined palettes are carefully curated by experts, often based on principles of color theory and perceptual research. A prominent example is ColorBrewer (Harrower2003ColorBrewerorgAO), a widely adopted tool that provides palettes for encoding sequential (ordered values), diverging (centered around a critical midpoint), and qualitative (categorical classes) data. These palettes are designed to maximize readability across print and digital media, and they have become a de facto standard in visualization practice. Similarly, visualization platforms such as Tableau provide pre-defined color palettes embedded into their systems (Milligan2018Tableau1C), allowing designers to quickly apply professionally designed schemes. While effective for common tasks, pre-defined palettes are inherently limited in flexibility. They typically support only a small set of colors and do not adapt to the specific characteristics of new datasets or visualization contexts, making them less suitable for complex visualizations requiring larger or interdependent color sets.
Automatic palette generation methods attempt to overcome these limitations by producing custom palettes that adapt to particular visualization needs. These methods often employ optimization algorithms or learning-based approaches to balance perceptual discriminability and aesthetic preference. For example, Colorgorical (Gramazio2017ColorgoricalCD) generates palettes by optimizing a user-defined balance between color discriminability and aesthetic preference for categorical data. The Palettailor (Lu2020PalettailorDC) incorporates data characteristics and employs a simulated annealing–based optimization when generating color palettes, assigning colors in a way that enhances visual discrimination between classes while preserving overall perceptual harmony. While such approaches focus on generating palettes for categorical data, our method tackles the cross-mode adaptation problem: transforming palettes designed for light mode into ones suitable for dark mode. To achieve this, we define the optimization problem around three dark mode–specific criteria, luminance contrast, semantic consistency, and relative color differences, ensuring that adapted palettes remain legible, meaningful, and visually coherent.
@@ 3.1. Dark Mode Visualization Transformation Factors
While there is extensive research on visualization design principles and color theory (schanda2007colorimetry; Healey1996ChoosingEC; Yuan2021InfoColorizerIR; Wang2019OptimizingCA), and major platforms provide dark mode UI guidelines (appleDarkMode; androidDarkMode), adapting visualizations specifically for dark mode remains unexplored. We analyzed dark mode UI design principles and identified three critical design factors as design requirements (DR) for effective dark mode visualization transformation:
  • DR1: Perceptual Discriminability. Visualization elements must remain clearly distinguishable against the dark background, just as they are in the light mode. Vision research further demonstrates that when luminance contrast decreases, color differences play a critical role in preserving discriminability (Chen2022ColorPolarity). This aligns with industry guidelines for dark mode from Android (androidDarkMode) and iOS (appleDarkMode), and with visualization research emphasizing perceptual separability through optimized color assignment (Wang2019OptimizingCA).
  • DR2: Color Semantic Consistency. Colors must maintain their semantic relationships between the light and the dark modes (Erickson2020DarkLightMA; Andrew2024LightAD). For example, in visualization contexts where red indicates negative values and green indicates positive values, these color-meaning associations should remain consistent across modes. Schloss et al. (Schloss2019ColorMeaning) show that colormaps are effective when color–quantity mappings align with users’ expectations (e.g., darker colors map to larger quantities). Preserving these semantic relationships in dark mode is therefore essential to ensure consistency and interpretability.
  • DR3: Relative Perceptual Differences. The relative visual differences between adjacent colors must be preserved when transforming to dark mode. For example, in a visualization where highlighted data points (e.g., outliers) are designed to stand out from regular data points through color contrast, this visual emphasis should remain equally noticeable after transformation (Lu2023InteractiveCC). If a bright orange highlight distinctly contrasts with blue regular points in light mode, their transformed colors in dark mode should maintain this level of differentiation. Preserving this relative color differences ensure that visual hierarchy and emphasis techniques remain effective across modes (Mittelstdt2014MethodsFC).
@@ 3.3. Color Palette Extraction
To extract the color palette from the input image, Chameleon uses k-means clustering (Hartigan1979AKC) on the RGB values of all non-background pixels, where each pixel is treated as a three-dimensional point (R, G, B). The clustering process groups similar colored pixels together and identifies $k$ cluster centroids, which form our extracted color palette. As shown in Figure 2, the choice of $k$ affects the granularity of color identification, where a larger $k$ value captures more distinct colors and subtle variations in the visualization. For example, when kk=5, the algorithm identifies only the most dominant colors, while kk=15 captures more nuanced color variations. For continuous visualizations (Figure 3), such as heatmaps or density plots, the impact of $k$ becomes even more significant. A higher number of clusters ensures smoother transitions between colors, as demonstrated in Figure 3, where kk=30 produces smoother gradients than kk=10. This flexibility in $k$ selection allows Chameleon to handle both visualizations with distinct color categories and those requiring smooth color transitions.
@@ 3.4. Loss Functions and Design
Luminance Contrast Consistency: Luminance contrast plays a critical role in object-background differentiation (Zhang2023DontPA; Lubin1997AHV). A sufficient level of luminance contrast enables the human eye to distinguish the visual elements from the background, thereby ensuring the clarity of the visualization. In our work, the function aims to maintain the luminance contrast between visualization elements with the dark mode background is comparable to that of the light mode background. This allows the human eye to perceive the same contrast of elements in the dark mode background as over the light mode background, thereby maintaining clarity across different modes. In our approach, we utilize the L channel from the LCH color space (schanda2007colorimetry), as it closely matches human perception of luminance (Hanbury2011MATHEMATICALMI).
@@
Color Consistency: In many visualizations, color choices carry semantic meaning (e.g.,  Figure 1). Maintaining color consistency helps users to interpret the data in both light and dark modes without confusion. This function aims to preserve the semantic meaning relationships of colors by ensuring that the colors in the dark mode remain perceptually similar to those in the light mode. We evaluate color consistency using the Delta CIE2000 metric (Sharma2005TheCC), which measures perceptual color differences:
@@
where $C_{i}^{\text{light}}$ represents a specific color in the original light mode visualization, and $C_{i}^{\text{dark}}$ represents its corresponding transformed color in the dark mode. $\Delta E_{2000}$ represents the color difference between the light mode and dark mode color $C_{i}$. The $\Delta E_{2000}$ metric was chosen because it better matches human perception of color difference compared to earlier color difference formulas like CIELAB ($\Delta E_{ab}$) or Delta CIE94 ($\Delta E_{94}$(Sharma2005TheCC). This formulation ensures that transformed colors remain perceptually close to their originals, preserving semantic color meaning across modes (DR2).

### v2

@@ Abstract.
Figure 1. Visualization examples adapted from Shen et al. (Shen et al., 2017), demonstrating four different rendering conditions. (A) Original visualization in light mode. (B) A typical dark mode conversion, where only the background is darkened while the original foreground colors are left unchanged, often leading to poor contrast and reduced legibility. (C) Visualization rendered using direct color inversion, which can distort the intended color semantics and lead to misinterpretation. (D) Visualization transformed using our method, Chameleon, which optimizes color contrast and preserves semantic color mappings for improved legibility in dark mode.
@@ 1. Introduction
In recent years, dark mode has gained popularity across mobile platforms like Android and iOS due to its benefits, such as reduced eye strain in low-light conditions and prolonged battery life on OLED and AMOLED screens (Andrew et al., 2024; Iyer et al., 2003). Dark mode is a user interface (UI) theme that uses a dark background with light text, as opposed to the default light mode, which typically uses a light background with dark text. Users can easily switch to dark mode through system settings, resulting in an immediate change in the system’s UI. Visualizations are commonly used on mobile devices because of their convenience (Choe et al., 2019; Lee et al., 2020; Roberts et al., 2014). When dark mode visualizations are effectively designed, they minimize distractions (Apple, 2025b), make important content stand out, and reduce eye strain (Qiao and Wu, 2023; Popelka et al., 2022). They can also be made to adjust to changing light conditions, improving user comfort in dim environments (Qiao and Wu, 2023). For instance, Mairena et al. (Mairena et al., 2022) found that the users performed better using the dark mode scatterplots, with faster search time and higher accuracy, compared to light mode.
@@
However, most visualizations were designed for light mode, and when the system switches to dark mode, the mismatch between visualization elements and the dark background creates significant visibility problems. Figure 1[Uncaptioned image] shows a visualization from a scientific paper (Shen et al., 2017) that is clearly visible against a light background. Yet, when the system switches to dark mode, as shown in Figure 1[Uncaptioned image], the visualization elements become difficult to discern against the dark background, compromising their legibility. A common shortcut is to simply invert the color palette (Andrew et al., 2024), as shown in Figure 1[Uncaptioned image]. Unfortunately, this inversion cannot preserve the color semantic relationships between the inverted visualization and the original visualization: the blue color in the original visualization is inverted to yellow, and the orange color is inverted to blue. This change largely disrupts users’ cognitive ability to recognize and interpret the visualization across both modes.
@@
As a consequence, designers are left to manually adapt visualizations for dark mode. This approach requires designers to build dark mode visualizations from scratch, a time-consuming process that involves complex multi-factor considerations: maintaining adequate contrast against the dark background, preserving the original color relationships, and ensuring overall visual harmony. These factors often conflict with each other. For example, increasing contrast for better legibility might disrupt color relationships or visual harmony. Since such factors are interdependent, designers must engage in extensive trial and error, making repeated adjustments to balance these competing requirements. While operating systems like Android and iOS offer guidelines (Apple, 2025a, b; Material Design (2025), Google; Pinterest, 2025) for UI design in dark mode, these guidelines typically focus on discrete UI elements like buttons and menus. As a consequence, the guidelines are often too abstract to be directly applicable to the specific needs of visualizations (Andrew et al., 2024; Qiao and Wu, 2023), which requires more nuanced design considerations.
@@
To reduce designers’ workload, there is a need for an automated tool that can transform light mode visualizations to dark mode by adjusting their color palettes. While there are existing tools for automatic color palette generation (Lu et al., 2021; Gramazio et al., 2017; Nardini et al., 2021), these cannot be directly applied to dark mode transformation. These tools focus solely on generating optimized palettes for single mode and do not address the challenges of cross-mode transformation. These single-mode approaches creates several problems. First, colors that appear legible in light mode may become harsh or difficult to distinguish against dark backgrounds, an issue not considered by existing automated methods. Second, colors need to maintain consistency between the modes to help users quickly interpret visualizations regardless of display mode. Finally, visual harmony and hierarchy must be preserved. For example, if highlighted regions have distinct contrast with adjacent areas in the light mode, this relationship should persist in the dark mode for consistent information emphasis. These cross-mode relationships and dark background interactions present unique challenges that require a different approach to the color palette generation.
@@
To fill this gap, we propose Chameleon, an algorithm designed to automatically transform visualizations from light mode to dark mode. Through literature reviews (Apple, 2025b; Material Design (2025), Google; Lu et al., 2021; Gramazio et al., 2017; Nardini et al., 2021), we identified three critical factors in dark mode visualization design: maintaining adequate luminance contrast for legibility, preserving color semantic relationships, and ensuring visual harmony between adjacent colors. The key challenge lies in quantifying these factors and balancing them effectively. Chameleon addresses this by formulating the transformation as an optimization problem that mathematically models these design factors. Specifically, Chameleon takes an image of a light mode visualization as input and analyzes its color palette. The algorithm then optimizes the dark mode color palette by balancing three designed loss functions: luminance contrast consistency to ensure legibility against the dark background, color consistency to maintain semantic relationships from the original visualization, and adjacent color difference consistency to preserve visual harmony and hierarchy. This automated optimization eliminates the initial trial-and-error phase, providing designers with an optimized starting point that they can fine-tune rather than starting from scratch. The weights of these loss functions can be adjusted if designers want to prioritize different factors according to their specific needs. As shown in Figure 1[Uncaptioned image] and Figure 1[Uncaptioned image], this approach produces a dark mode visualization that successfully balances visual clarity, color consistency, and visual harmony.
@@ 2.1. Color Palette Design Principles
Color palette design for charts centers on two main considerations: perceptual discriminability and aesthetics (Yuan et al., 2022). The perceptual discriminability ensures that users can easily distinguish between colors within a visualization (Gramazio et al., 2017; Stone et al., 2014), a requirement that is critical for accurate data interpretation. Healey (Healey, 1996) emphasized that “colors should be well-separated,” underscoring the importance of maximizing visual distinctness. Several factors influence discriminability: the data properties being represented (e.g., continuous vs. categorical data) (Levkowitz and Herman, 1992; Tennekes and de Jonge, 2014), the analytic tasks users perform (e.g., comparing classes, identifying outliers) (Tominski et al., 2008), the ease of color naming and categorical labeling (Heer and Stone, 2012), and the degree of perceptual contrast among colors (Wang et al., 2019; Mittelstädt et al., 2014). Healey (Healey, 1996) proposed selecting representative colors from ten hue regions in Munsell’s color space while maximizing perceptual distances between them, while Maxwell (Maxwell, 2000) similarly sought to maximize class discriminability, accounting for spatial distribution of colors. However, these approaches face limitations like the resulting palettes are not always aesthetically pleasing (Lu et al., 2021), stressing that a proper balance between perceptual discriminability and aesthetics.
Beyond discriminability, palettes must also be visually appealing, as aesthetic qualities influence user engagement and interpretability. Research shows that color harmony, the balanced arrangement of colors, enhances visual appeal and reduces cognitive strain (Meier et al., 2004). Furthermore, manipulating hue, saturation, and brightness allows designers to balance vividness with readability, ensuring that palettes remain both expressive and coherent (Moreland, 2009). Well-chosen palettes thus balance clarity and beauty, encouraging users to focus on insights rather than being distracted by jarring color choices. These design principles take on added importance in dark mode settings, where background inversion can drastically alter perceived contrast and harmony. Guidelines from major platforms such as Android (Material Design (2025), Google) and iOS (Apple, 2025b) stress the importance of maintaining strong luminance contrast between visualization elements and background. Without such care, a palette optimized for light mode may become unreadable, misleading, or aesthetically unbalanced in dark mode environments. In this work, we build on these principles to address the unique challenges of dark mode visualization.
@@ 2.2. Color Palette Generation
The existing approaches to color palette generation fall into two categories: pre-defined and automatic. Pre-defined palettes are carefully curated by experts, often based on principles of color theory and perceptual research. A prominent example is ColorBrewer (Harrower and Brewer, 2003), a widely adopted tool that provides palettes for encoding sequential (ordered values), diverging (centered around a critical midpoint), and qualitative (categorical classes) data. These palettes are designed to maximize readability across print and digital media, and they have become a de facto standard in visualization practice. Similarly, visualization platforms such as Tableau provide pre-defined color palettes embedded into their systems (Milligan and Guillevin, 2018), allowing designers to quickly apply professionally designed schemes. While effective for common tasks, pre-defined palettes are inherently limited in flexibility. They typically support only a small set of colors and do not adapt to the specific characteristics of new datasets or visualization contexts, making them less suitable for complex visualizations requiring larger or interdependent color sets.
Automatic palette generation methods attempt to overcome these limitations by producing custom palettes that adapt to particular visualization needs. These methods often employ optimization algorithms or learning-based approaches to balance perceptual discriminability and aesthetic preference. For example, Colorgorical (Gramazio et al., 2017) generates palettes by optimizing a user-defined balance between color discriminability and aesthetic preference for categorical data. The Palettailor (Lu et al., 2021) incorporates data characteristics and employs a simulated annealing–based optimization when generating color palettes, assigning colors in a way that enhances visual discrimination between classes while preserving overall perceptual harmony. While such approaches focus on generating palettes for categorical data, our method tackles the cross-mode adaptation problem: transforming palettes designed for light mode into ones suitable for dark mode. To achieve this, we define the optimization problem around three dark mode–specific criteria, luminance contrast, semantic consistency, and relative color differences, ensuring that adapted palettes remain legible, meaningful, and visually coherent.
@@ 3.1. Dark Mode Visualization Transformation Factors
While there is extensive research on visualization design principles and color theory (Schanda, 2007; Healey, 1996; Yuan et al., 2022; Wang et al., 2019), and major platforms provide dark mode UI guidelines (Apple, 2025b; Material Design (2025), Google), adapting visualizations specifically for dark mode remains unexplored. We analyzed dark mode UI design principles and identified three critical design factors as design requirements (DR) for effective dark mode visualization transformation:
  • DR1: Perceptual Discriminability. Visualization elements must remain clearly distinguishable against the dark background, just as they are in the light mode. Vision research further demonstrates that when luminance contrast decreases, color differences play a critical role in preserving discriminability (Chen and Muhamad, 2022). This aligns with industry guidelines for dark mode from Android (Material Design (2025), Google) and iOS (Apple, 2025b), and with visualization research emphasizing perceptual separability through optimized color assignment (Wang et al., 2019).
  • DR2: Color Semantic Consistency. Colors must maintain their semantic relationships between the light and the dark modes (Erickson et al., 2020; Andrew et al., 2024). For example, in visualization contexts where red indicates negative values and green indicates positive values, these color-meaning associations should remain consistent across modes. Schloss et al. (Schloss et al., 2019) show that colormaps are effective when color–quantity mappings align with users’ expectations (e.g., darker colors map to larger quantities). Preserving these semantic relationships in dark mode is therefore essential to ensure consistency and interpretability.
  • DR3: Relative Perceptual Differences. The relative visual differences between adjacent colors must be preserved when transforming to dark mode. For example, in a visualization where highlighted data points (e.g., outliers) are designed to stand out from regular data points through color contrast, this visual emphasis should remain equally noticeable after transformation (Lu et al., 2023). If a bright orange highlight distinctly contrasts with blue regular points in light mode, their transformed colors in dark mode should maintain this level of differentiation. Preserving this relative color differences ensure that visual hierarchy and emphasis techniques remain effective across modes (Mittelstädt et al., 2014).
@@ 3.3. Color Palette Extraction
To extract the color palette from the input image, Chameleon uses k-means clustering (Hartigan and Wong, 1979) on the RGB values of all non-background pixels, where each pixel is treated as a three-dimensional point (R, G, B). The clustering process groups similar colored pixels together and identifies $k$ cluster centroids, which form our extracted color palette. As shown in Figure 2, the choice of $k$ affects the granularity of color identification, where a larger $k$ value captures more distinct colors and subtle variations in the visualization. For example, when kk=5, the algorithm identifies only the most dominant colors, while kk=15 captures more nuanced color variations. For continuous visualizations (Figure 3), such as heatmaps or density plots, the impact of $k$ becomes even more significant. A higher number of clusters ensures smoother transitions between colors, as demonstrated in Figure 3, where kk=30 produces smoother gradients than kk=10. This flexibility in $k$ selection allows Chameleon to handle both visualizations with distinct color categories and those requiring smooth color transitions.
@@ 3.4. Loss Functions and Design
Luminance Contrast Consistency: Luminance contrast plays a critical role in object-background differentiation (Zhang et al., 2023; Lubin, 1997). A sufficient level of luminance contrast enables the human eye to distinguish the visual elements from the background, thereby ensuring the clarity of the visualization. In our work, the function aims to maintain the luminance contrast between visualization elements with the dark mode background is comparable to that of the light mode background. This allows the human eye to perceive the same contrast of elements in the dark mode background as over the light mode background, thereby maintaining clarity across different modes. In our approach, we utilize the L channel from the LCH color space (Schanda, 2007), as it closely matches human perception of luminance (Hanbury and Serra, 2011).
@@
Color Consistency: In many visualizations, color choices carry semantic meaning (e.g.,  Figure 1). Maintaining color consistency helps users to interpret the data in both light and dark modes without confusion. This function aims to preserve the semantic meaning relationships of colors by ensuring that the colors in the dark mode remain perceptually similar to those in the light mode. We evaluate color consistency using the Delta CIE2000 metric (Sharma et al., 2005), which measures perceptual color differences:
@@
where $C_{i}^{\text{light}}$ represents a specific color in the original light mode visualization, and $C_{i}^{\text{dark}}$ represents its corresponding transformed color in the dark mode. $\Delta E_{2000}$ represents the color difference between the light mode and dark mode color $C_{i}$. The $\Delta E_{2000}$ metric was chosen because it better matches human perception of color difference compared to earlier color difference formulas like CIELAB ($\Delta E_{ab}$) or Delta CIE94 ($\Delta E_{94}$(Sharma et al., 2005). This formulation ensures that transformed colors remain perceptually close to their originals, preserving semantic color meaning across modes (DR2).